### v1

# Multi-lingual Functional Evaluation for Large Language Models

Abstract
Multi-lingual competence in large language models is often evaluated via static data benchmarks such as Belebele, M-MMLU and M-GSM. However, these evaluations often fail to provide an adequate understanding of the practical performance and robustness of models across multi-lingual settings. In response, we create multi-lingual functional benchmarks – Cross-Lingual Grade School Math Symbolic (CL-GSM Symbolic) and Cross-Lingual Instruction-Following Eval (CL-IFEval)– by translating existing functional benchmark templates from English to five additional languages that span the range of resources available for NLP: French, Spanish, Hindi, Arabic and Yoruba. Our results reveal that some static multi-lingual benchmarks capture functional performance much more closely than others (i.e. across models, there is a 24%, 17% and 18% decrease in performance between M-GSM and CL-GSM Symbolic in English, French and Spanish respectively; similarly there’s a 15 - 24% performance drop across languages between Belebele and CL-IFEval, and only a 0.5% to 3% performance drop between M-MMLU and CL-IFEval). Similarly, we find that model robustness across languages varies significantly, with certain languages (eg. Arabic, English) being the most consistently well performing across evaluation iterations.
Multi-lingual Functional Evaluation for Large Language Models
Victor Ojewale1,  Inioluwa Deborah Raji2,  Suresh Venkatasubramanian1 1The Center for Tech Responsibility, Brown University, USA 2University of California, Berkeley, USA
1 Introduction
Figure 1: Description of the functional evaluation paradigm. Unlike with static data benchmarks, in the functional evaluation paradigm, model input prompts are not fixed but generated through a fixed template and a set of variables $X$ (modifiable prompt attributes meant to impact model outputs) and a set of distractors $D$ (modifiable prompt attributes meant to be ignored). The ground truth in this setting is generated through a fixed functional transformation $f(X)$. For instance, the prompt "Sally bought 2 red apples and 3 green apples. How much fruit did Sally buy?" is generated from the fixed template "$\{name\}$ bought $\{n_{1}\}$$\{color_{1}\}$ apples and $\{n_{2}\}$$\{color_{2}\}$ apples. How much fruit did $\{name\}$ buy?". This template involves the variables $X=\{n_{1},n_{2}\}$ and the distractors $D=\{name,color_{1},color_{2}\}$. The correct fixed output function in this case is $f(X)=n_{1}+n_{2}$.
Despite some meaningful progress, models operating in languages other than English have been regularly found to be more biased Talat et al. (2022), less safe Yong et al. (2023) and overall meaningfully less performant and robust  Ojo et al. (2023).
Popular multi-lingual LLM evaluations, such as Multilingual MMLU (M-MMLU) and the Multilingual Grade School Math (M-GSM) benchmark(Hendrycks et al., 2021; Shi et al., 2022; Cobbe et al., 2021), while useful, often fail to capture more meaningful indications of functional multi-lingual model performance – that is, the robust execution of a given prompt across a variety of languages (see Figure  1). In this paper, we extend the scope of two English "functional" evaluation datasets – IFEval Zhou et al. (2023), and GSM-Symbolic Mirzadeh et al. (2024) – by translating its prompt templates into five additional languages: French, Spanish, Hindi, Arabic and Yoruba.
Our experiments reveal that there are significant, though inconsistent disparities between multi-lingual model performance on functional vs static data benchmarks – in the vast majority of cases, the models perform much better on the static data benchmarks than the functional evaluations, though these differences are much larger for certain static data benchmarks than others. The performance gap between languages is comparable in both cases but this varies widely across languages and models. Furthermore, functional benchmark results reveal robustness inconsistencies across languages, with models performing more robustly in certain languages over others for specific prompt templates or question types.
Refer to captionRefer to captionRefer to caption
Figure 2: Correlation plot of Performance Gap between MGSM, MMLU, Belebele (left-to-right) and and CL-IFEval for High Resourced Languages only (en, fr, es). This reveals that measured language performance gaps (i.e. the difference between the performance on the highest performant language and the lowest performant language) are notably larger in functional evaluations than in static data benchmarks.
2 Related Work
Common multi-lingual data benchmarks such as M-MMLU (Hendrycks et al., 2021; Lai et al., 2023a), FLORES (Goyal et al., 2022), BeleBele (Bandarkar et al., 2024), and XLSum (Hasan et al., 2021) possess known limitations. The use of direct translation for many of these benchmarks has been critiqued by some as being devoid of realistic cultural context Romanou et al. (2024); Singh et al. (2024). Furthermore, research reveals that English language benchmark data contamination might distort reported benchmark performance in English or possibly additional languages (as is the case for MMLU Dodge et al. (2021) and GSM Zhang et al. (2024)).
Functional evaluation involves "templating" a common popular benchmark with modifiable variables. For example, GSM-Symbolic templates examples of the static data benchmark GSM8k  Cobbe et al. (2021) to generate the input permutations. The ground truth for the math problems is then calculated using literal template-based functional mappings from input values to the expected output (see Figure  1). Recent work has attempted to set up similar symbolic annotations for natural language benchmarks Hennigen et al. (2023), and we can see a similar verifiable, function-based template format with instruction-following benchmarks (Liu et al., 2024; Chang et al., 2023) such as the IFEval datasetZhou et al. (2023). Although some concurrent work – a proposed Multi-lingual IFEval (M-IFEval)  Dussolle et al. (2025) – has attempted to translate the IFEval template to Spanish, French and Japanese, we have yet to see more systematic analysis of how such functional evaluations can inform better assessments of performance and robustness in multi-lingual deployment settings.
Benchmark Avail. Lang Prompts/Lang Templates Samples/Template Task
Functional Benchmarks
Cross-Lingual GSM Symbolic (Ours) 5 5000 100 50 Math Reasoning
Cross-Lingual IFEval (Ours) 5 541 10 31 - 163 Instruction following
Static Benchmarks
Multilingual MMLU Lai et al. (2023a) 26 13062 N/A N/A MC knowledge recall
Multilingual GSM Shi et al. (2022) 10 250 N/A N/A Math Reasoning
Belebele Bandarkar et al. (2024) 122 900 N/A N/A MC reading comprehension
Table 1: Descriptive Statistics of Analyzed Benchmarks.
3 Benchmark Datasets and Templates
To make multi-lingual functional benchmarks, we translate 541 English IFEval prompts, each containing at least one verifiable instruction, as well as 5000 question-answer pairs generated from 100 templated symbolic problem variants in the English GSM-Symbolic dataset, into French, Spanish, Yoruba, and Hindi. In line with Bang et al. (2023); Lai et al. (2023b), we refer to languages by their ISO 639-1 language code abbreviations and classify languages by the data ratio of the representation of that language in CommonCrawl (see 3) – these languages were thus selected to cover a spectrum of high-resource and lower-resource contexts. As with prior work Montariol et al. (2022); Lai et al. (2023a); Chen et al. (2023), translations are conducted with Google Translate (Wu et al., 2016). To confirm translation quality, we conduct spot checks (i.e. the manual analysis of 10-20 examples per datasets) with native and proficient language speakers to identify inconsistencies in translation accuracy. We name the resulting functional evaluation datasets Cross-Lingual IFEval (CL-IFEval)and Cross-Lingual GSM Symbolic (CL-GSMSym). All translated prompts can be found in our uploaded supplementary materials. For comparison with a set of multi-lingual static data benchmarks, we use Multilingual MMLU (M-MMLU) Lai et al. (2023a); Hendrycks et al. (2021), Multilingual Grade School Math benchmark (MGSM) Shi et al. (2022), and the Belebele benchmark  Bandarkar et al. (2024). The Cross-Lingual GSM Symbolic and Cross-Lingual IFEval datasets are publicly available.  [^1] https://huggingface.co/datasets/vojewale/Cross-lingualGSMSymbolic [^2] https://huggingface.co/datasets/vojewale/Cross-lingualIFEval
Further dataset details can be found in Table  1.
4 Model Evaluation
For reproducible consistency in reporting, our evaluations spanned open-source large language models, including instruction-tuned and multilingual variants - Aya 23-35B, Aya Expanse-32B, Gemma-2-9B-it, Qwen3-8B, Mistral-7B-Instruct-v0.3, Mixtral-8x7B-Instruct-v0.1. For further comparison on CL-IFEval, we also evaluate on the proprietary models GPT-4o-mini and Claude Sonnet 3.5. Using CL-IFEval, we calculate the prompt-level strict/loose accuracy and the instruction-level strict/loose accuracy as defined in the original IFEval paper Zhou et al. (2023). For CL-GSMSym, we evaluate a random sample of 500 translated question-answer pairs per language using the language-specific 8-shot setting for GSM8K evaluation.
5 Results
Figure 3: Cross-Lingual IFEval Strict Prompt Accuracy.
Figure 4: Cross-Lingual IFEval Aya-23-35B Model Comparison Across Languages
Figure 5: Cross-Lingual GSMSym Aya-23-35B Template Comparison Across Languages
5.1 Average Performance Across Languages & Model Ranking
Models perform consistently high for high-resourced languages on static natural language benchmarks (eg. For M-MMLU, English (66% - 69%), French (51% - 65%) and Spanish (51% - 65%) performance is consistently higher; for Belebele, performance across models for high-resource languages is even stronger (79% – 90%)). However, there is notable performance variation across models for mathematical reasoning as measured with static data benchmarks – for example, M-GSM English scores range from 44% to 86% across models, while performance in French and Spanish is generally lower, though still relatively competitive and wide-ranging (French: 44% - 76%, Spanish: 37% - 84%). When available, performance across models decreases on medium resource languages - with scores ranging from 31% to 48% in Hindi for M-MMLU, and 41% to 74% in Hindi for Belebele. Low-resource languages like Yoruba show drastic performance drops, with scores below 32% on Belebele.
On the other hand, functional benchmarks in natural language, such as CL-IFEval, show a wide range on even English performance (46% - 81%), with other high and medium resource languages such as French (31%- 58%), Spanish (29% - 54%), and Arabic (20% - 49%) being similarly wide ranging in performance. CL-GSMSym shows a similar pattern, with a wide range of model performance in high and medium resource languages (English: 49% - 87%; French: 39% - 78%; Spanish: 34% - 82%; Arabic: 18% - 79%; Hindi: 15% - 72%). Interestingly, in both cases, performance for low resource languages was more consistently low (e.g. in Yoruba, performance on CL-IFEval ranges from 9% to 16%, and performance on CL-GSMSym ranges from 4% to 14%).
For both static benchmarks, the same models are consistently the most performant - typically Qwen3-8B, followed by either Gemma-2-9b-it or Aya-expanse-32B - with the other models not far behind. For functional benchmarks there is a clear discrepancy between a performant class of models (Qwen3-8B, followed by Aya-expanse-32B, followed by Gemma-2-9b-it) and the rest of the models. Unlike with static data benchmarks, this stark performance discrepancy persists, even for low resource languages. For example, on CL-IFEval, the top three models have an English performance range of 70% to 82%, whereas the bottom three models all have an English performance range consistently around 46%. In Yoruba, this performance class discrepancy is still observed, though the performance gap between models is much more narrow. Comparatively, the top three models on the static data benchmark Belebele perform in English at 90% to 93% and the bottom three models perform at 80% to 86%. Across languages, we can see model rankings change significantly when assessed with static vs functional benchmarking. For example, a typically lower-ranked model like Aya-23-35B, outranks Aya-expanse-32B and Qwen3-8B in Yoruba on the static Belebele benchmark but not in Yoruba on the functional CL-IFEval benchmark. Similarly, the Aya-23-35B model outranks Gemma-2-9b-it generally on the static M-GSM, even though it consistently performs worse than Gemma-2-9b-it on the functional CL-GSMSym benchmark, for the same languages.
Full details of the performance results for the static data benchmarks and functional benchmarks can be found in Appendix D and E.
5.2 Language Performance Gap
We define the language performance gap to be the difference of the model’s accuracy on its lowest performant language, and the highest performant language. For high- resourced languages, functional evaluations tend to reveal a much larger performance discrepancy than static data benchmarks (see Figure  2). Notably, Aya-expanse-32B, a Cohere model marketed specifically for its multi-lingual capability, has a 5.89% average error gap between high-resourced languages (en, es, fr) across static data benchmarks, but a 23.47% performance gap between those same languages on the functional benchmark CL-IFEval. Similarly, Qwern3-8b has fairly low language performance gaps across high resourced languages on static data benchmarks like M-GSM (14.4%), M-MMLU (4.78%), and Belebele (3.00%), but a large language performance gap on the functional CL-IFEval benchmark (27.36%). Interestingly, for mathematical reasoning, the opposite can be true – for example, the language performance gap across high resource languages is lower for the functional CL-GSMSym (e.g. for Qwern3-8b, 8.6%) than the static data benchmark M-GSM (eg. for Qwern3-8b, 14.4%).
Similarly, when including low resource language settings, CL-IFEval tends to show more optimistic numbers (a higher average performance, a lower performance gap) than the static data benchmark, Belebele. In Appendix B, Tables 4, 5,6, as well as Figure  2, we provide further details on these results.
5.3 Instruction and Template-Level Robustness Across Languages
To evaluate instruction-level robustness across languages, we use CL-IFEval to compare model behavior on grouped instruction categories. Figures 6, 7, and 8 show success rates for Aya-23-35B, Gemma-2-9B, and Qwen3-8B respectively. These plots reveal notable variation in cross-lingual generalization, with Yoruba performing the worst across most instruction groups and showing complete failure in the “Start/End” category for all models.
To further probe model robustness under controlled prompt variations, we use a subset of CL-GSMSym consisting of 50 samples generated from a fixed set of 10 mathematical question templates. Figures 9, 10, and 11 display the per-template accuracy of the same three models across languages. Template 3(Appendix H), which involves probabilistic inference, consistently produced the lowest performance across models and languages highlighting a persistent weakness in generalizing probabilistic reasoning across linguistic contexts. Interestingly, performance in certain middle-resourced languages (ar, hi) is just as if not more robust for certain types of problems than high resourced languages (e.g. in CL-GSM Symbolic, see Template 2, 4 and 10 results in Figure  9).
6 Conclusion
We introduce two new benchmarks, CL-IFEval and CL-GSMSym, for multi-lingual functional evaluation. Our experiments uncovered major language performance gaps across languages, even for LLMs with robust multilingual claims and strong static data benchmark scores.
Limitations.
We use automated translation tools like Google Translate in constructing CL-IFEval and CL-GSMSym. While these tools offer broad language coverage and facilitate large-scale data generation, they introduce potential inaccuracies, particularly for lower resourced languages like Yoruba, and when dealing with conversions across metric and imperial measurement systems.
Another limitation is the emphasis in our analysis on open-weight models. While we evaluate gpt-4o-mini and Claude Sonnet 3.5 with CL-IFEval as a commercial baseline, our primary focus remains on open-weight models such as Mixtral-8x7B, Mistral-7B, Gemma2-9B-it, Qwen3-8B and AYA models. This creates a possible inherent bias in the scope of our comparisons, as proprietary models may perform significantly better than their open-weight counter-parts. On the other hand, the consistency and transparency of open-weight models make them the preferable object of study – the incorporation of proprietary models can make results hard to reproduce reliably.
Future Work.
Future directions include systematically curating higher-quality translations, expanding into multi-lingual code or multi-modal instruction evaluation, and further investigating the robustness and error patterns in both functional and static benchmarks. Also, as functional evaluations involve automatic verification, there is some possibility of extrapolating this framework in the training and evaluation of multi-lingual reasoning models Yong et al. (2025).
Acknowledgments
The authors would like to thank Zheng-Xin Yong for feedback on the work. This work was supported in part by the MacArthur Foundation, the Mozilla Foundation, and the Heising-Simons Foundation.
Appendix A Additional Experimental Details
Benchmark Avail. Lang Prompts/Lang Templates Samples/Template Task
Functional Benchmarks
Cross-Lingual GSM Symbolic (Ours) 5 5000 100 50 Math Reasoning
Cross-Lingual IFEval (Ours) 5 541 10 31 - 163 Instruction following
Static Benchmarks
Multilingual MMLU Lai et al. (2023a) 26 13062 N/A N/A MC knowledge recall
Multilingual GSM Shi et al. (2022) 10 250 N/A N/A Math Reasoning
Belebele Bandarkar et al. (2024) 122 900 N/A N/A MC reading comprehension
Table 2: Descriptive Statistics of Analyzed Benchmarks.
A.1 Models
Our evaluations spanned open-source large language models, including instruction-tuned and multilingual variants. Below, we provide details on each model:
  • Aya 23-35B: A multilingual instruction-tuned model based on Cohere’s Command framework (Aryabumi et al., 2024).
  • Aya Expanse-32B: Part of the Aya Expanse series, designed to enhance multilingual performance capabilities(Dang et al., 2024).
  • Gemma-2-9B-it: An instruction-tuned model trained on 8 trillion tokens from diverse sources, including web documents, code, and mathematical text. It is optimized for a wide range of text generation tasks, including question answering, summarization, and code understanding (Team et al., 2024).
  • Qwen3-8B: A dense and mixture-of-experts (MoE) model supporting 100+ languages. It includes specialised modes for logical reasoning, code generation, and agent-based tasks (Yang et al., 2025).
  • Mistral-7B-Instruct-v0.3: An open-source instruction fine-tuned variant of Mistral-7B (Jiang et al., 2023).
  • Mixtral-8x7B-Instruct-v0.1: (Jiang et al., 2024)A sparse mixture-of-experts model with 12.9B active parameters per token, fine-tuned for instruction tasks.
For comparison on CL-IFEval, we also evaluate on the proprietary models GPT-4o-mini and Claude Sonnet 3.5.
A.2 Language Resource Classification
Table 3 classifies the languages used in our evaluation based on their relative resource levels in CommonCrawl.
Language CC-MAIN-2025-05 (%) Category
English (en) 43.37 High-resource (HRL)
French (fr) 4.52 High-resource (HRL)
Spanish (es) 4.64 High-resource (HRL)
Arabic (ar) 0.66 Medium-resource (MRL)
Hindi (hi) 0.20 Medium-resource (MRL)
Yoruba (yo) 0.0009 Extremely low-resource (X-LRL)
Table 3: Classification of languages based on resource availability in CommonCrawl (CC-MAIN-2025-05).
Appendix B Full Performance Gap Results
B.1 High-Resource Language Performance Gaps
Model CL-IFEval CL-GSMSym M-GSM M-MMLU Belebele
Aya-23-35B 15.90 4.4 8.0 4.63 5.00
Mixtral-8x7B-Instruct 16.63 7.0 12.0 3.63 6.89
Mistral-7B-Instruct 12.39 14.0 8.0 7.69 15.88
Aya-expanse-32B 23.47 6.2 9.6 3.84 4.22
Gemma-2-9b-it 21.45 27.4 29.2 6.39 3.66
Qwen3-8b 27.36 8.6 14.4 4.78 3.00
Table 4: Performance Gap on High-Resource Languages, as measured by the benchmark accuracy difference between the results on the best performant language and worst performant language across the set of analyzed languages (en, fr, es).
B.2 High to Medium-Resource Language Performance Gaps
Model CL-IFEval CL-GSMSym M-MMLU Belebele
Aya-23-35B 15.90 17.6 18.08 22.45
Mixtral-8x7B-Instruct 26.06 33.2 30.65 36.78
Mistral-7B-Instruct 28.48 33.8 27.92 38.44
Aya-expanse-32B 28.84 11.0 15.29 15.45
Gemma-2-9b-it 27.18 27.4 18.20 19.55
Qwen3-8b 34.57 27.4 20.03 23.45
Table 5: Performance Gap on High to Medium-Resource Languages, as measured by the benchmark accuracy difference between the results on the best performant language and worst performant language across the set of analyzed languages (en, fr, es, ar, hi).
B.3 High to X Low-Resource Language Gaps
Model CL-IFEval CL-GSMSym Belebele
Aya-23-35B 33.09 57.4 54.12
Mixtral-8x7B-Instruct 36.41 56.8 57.11
Mistral-7B-Instruct 33.28 44.2 50.11
Aya-expanse-32B 56.38 71.6 60.45
Gemma-2-9b-it 53.42 64.4 52.44
Qwen3-8b 67.65 73.0 62.89
Table 6: Performance Gap on High to X Low-Resource Languages, as measured by the benchmark accuracy difference between the results on the best performant language and worst performant language across the set of analyzed languages (en, fr, es, ar, hi, yo).
Appendix C Static Data Benchmark Results
C.1 Multilingual MMLU (M-MMLU)
The M-MMLU benchmark assesses knowledge and reasoning in a wide range of subjects and languages. The results in Table 11 demonstrate that performance is generally higher in English (0.66–0.69) and other well-resourced languages such as French (0.51–0.65) and Spanish (0.51–0.65). In contrast, performance in medium-resource languages like Hindi is considerably lower, with scores ranging from 0.31 to 0.48. Arabic performance falls between these extremes but is still notably below English and French levels.
C.2 Multilingual Grade School Math benchmark (MGSM)
MGSM tests a model’s capacity to handle multi-step reasoning problems, particularly in mathematical contexts. The results in Table 12 reveal a stark performance gap between English and other languages. English scores range from 0.44 to 0.86, while performance in French and Spanish is generally lower, though still relatively competitive (French: 0.44–0.76, Spanish: 0.37–0.84). French and Spanish scores perform better than seen in M-MMLU, but medium and low-resource languages are not included in this benchmark.
C.3 Belebele Evaluation
The Belebele benchmark Bandarkar et al. (2024) assesses reading comprehension by requiring models to answer multiple-choice questions across a wide variety of languages. Table 10 highlights significant disparities in performance across resource levels. English comprehension is consistently strong, with scores around 0.85–0.90, while other high-resource languages such as French and Spanish achieve similarly high performance (0.79–0.88). However, low-resource languages like Yoruba show drastic performance drops, with scores below 0.32. Hindi also exhibits relatively weak results (0.41–0.74).
Appendix D Full Functional Benchmark Results
Model English French Spanish Arabic Hindi Yoruba
Aya-23-35B 46.21 31.60 30.31 35.86 36.23 13.12
Mixtral-8x7B-Instruct 45.84 31.42 29.21 25.32 19.78 9.43
Mistral-7B-Instruct 46.40 34.01 35.30 19.96 17.92 13.12
Aya-expanse-32B 70.24 50.83 46.77 44.18 41.40 13.86
Gemma-2-9b-it 69.87 51.94 48.42 45.1 42.69 16.45
Qwen3-8b 81.88 57.67 54.52 48.79 47.31 14.23
Table 7: Cross-Lingual-IFEval Prompt-level Strict performance across different models and languages.
Language Model Strict PL (%) Strict IL (%) Loose PL (%) Loose IL (%)
English GPT-4o-mini 81.14 86.81 84.47 89.32
AYA 23-35B 46.21 56.83 49.16 60.07
AYA Exp-32B 70.24 77.94 78.74 84.89
Claude Sonnet 85.76 90.04 89.64 92.92
Mistral 46.40 57.79 57.12 67.87
Mixtral 45.84 57.19 55.08 65.95
Gemma2-9B 69.87 78.77 73.75 82.01
Qwen3-8B 81.88 88.00 84.84 89.92
French GPT-4o-mini 55.08 66.42 58.22 69.42
AYA 23-35B 31.60 43.16 34.75 46.76
AYA Exp-32B 50.83 61.87 57.86 68.35
Claude Sonnet 62.84 72.42 63.58 73.50
Mistral 34.01 45.92 43.99 55.88
Mixtral 31.42 43.65 37.71 51.08
Gemma2-9B 51.94 62.94 53.41 64.74
Qwen3-8B 57.67 69.06 59.51 70.50
Spanish GPT-4o-mini 53.23 64.62 56.74 68.10
AYA 23-35B 30.31 42.20 33.82 45.56
Gemma 7B 30.13 41.13 34.57 46.88
AYA Exp-32B 46.77 58.51 53.23 65.35
Claude Sonnet 58.96 69.42 60.25 70.62
Mistral 35.30 46.16 37.70 48.92
Mixtral 29.21 40.53 36.41 49.28
Gemma2-9B 48.42 60.55 51.20 63.18
Qwen3-8B 54.52 66.30 56.56 68.22
Yoruba GPT-4o-mini 35.67 47.72 39.00 51.31
AYA 23-35B 13.12 23.50 14.23 24.58
AYA Exp-32B 13.86 23.86 20.33 32.61
Claude Sonnet 34.19 46.28 37.89 50.59
Mistral 13.12 24.22 17.56 30.94
Mixtral 9.43 19.06 13.86 25.90
Gemma2-9B 16.45 29.01 17.92 31.29
Qwen3-8B 14.23 25.17 15.71 27.21
Arabic GPT-4o-mini 47.87 61.15 50.27 63.42
AYA 23-35B 35.86 50.72 39.74 54.68
AYA Exp-32B 44.18 58.27 47.87 61.75
Claude Sonnet 53.23 65.10 54.15 65.94
Mistral 19.96 33.69 26.06 40.17
Mixtral 25.32 38.25 37.79 46.16
Gemma2-9B 45.10 57.79 46.39 59.35
Qwen3-8B 48.79 61.39 49.53 62.11
Hindi GPT-4o-mini 50.27 62.11 51.94 63.78
AYA 23-35B 36.23 47.48 40.48 53.12
AYA Exp-32B 41.40 53.12 46.95 59.11
Claude Sonnet 52.31 63.54 53.60 64.86
Mistral 17.92 29.25 24.21 37.77
Mixtral 19.78 31.18 26.25 38.97
Gemma2-9B 42.69 55.51 44.73 57.91
Qwen3-8B 47.31 59.35 49.53 61.75
Table 8: Comprehensive instruction-following results across multiple models and languages. PL = Prompt-Level Accuracy; IL = Instruction-Level Accuracy.
Model English French Spanish Arabic Hindi Yoruba
Aya-23-35B 56.00 55.00 59.40 61.20 43.60 3.80
Mixtral-8x7B-Instruct 56.80 55.00 62.00 32.40 28.80 5.20
Mistral-7B-Instruct 48.60 39.40 34.60 18.40 14.80 4.40
Aya-expanse-32B 82.60 76.40 82.20 76.80 71.60 11.00
Gemma-2-9b-it 77.00 49.60 54.80 55.20 57.00 12.60
Qwen3-8B 86.80 78.20 81.80 78.80 59.40 13.80
Table 9: CL-GSMSym (8-shot) performance across different models and languages.
Appendix E Full Static Data Benchmark Results
Model English French Spanish Arabic Hindi Yoruba
Aya-23-35B 0.8556 0.8133 0.8056 0.7667 0.6311 0.3144
Mixtral-8x7B-Instruct 0.8522 0.8100 0.7833 0.6344 0.4844 0.2811
Mistral-7B-Instruct 0.7944 0.6711 0.6356 0.5122 0.4100 0.2933
Aya-expanse-32B 0.8989 0.8844 0.8567 0.8367 0.7444 0.2944
Gemma-2-9b-it 0.9333 0.92 0.8967 0.8867 0.7378 0.4089
Qwen3-8B 0.9256 0.9089 0.8956 0.87 0.6911 0.2967
Table 10: Belebele performance across different models and languages.
Model English French Spanish Arabic Hindi
Aya-23-35B 0.6617 0.6168 0.6154 0.5394 0.4809
Mixtral-8x7B-Instruct 0.6854 0.6508 0.6491 0.4142 0.3789
Mistral-7B-Instruct 0.5900 0.5131 0.5181 0.3328 0.3108
Aya-expanse-32B 0.7396 0.7012 0.7025 0.6267 0.5867
Gemma-2-9b-it 0.7149 0.6635 0.6541 0.5565 0.5329
Qwen3-8B 0.7528 0.7050 0.7091 0.6072 0.5525
Table 11: M-MMLU(5-shot) performance across different models and languages.
Model English French Spanish
Aya-23-35B 0.668 0.588 0.600
Mixtral-8x7B-Instruct 0.632 0.512 0.536
Mistral-7B-Instruct 0.448 0.444 0.368
Aya-expanse-32B 0.856 0.760 0.840
Gemma-2-9b-it 0.664 0.372 0.512
Qwen3-8B 0.880 0.736 0.852
GPT-4o-mini 0.656 0.556 0.432
Claude 3.5 Sonnet 0.900 0.716 0.812
Table 12: MGSM (5-shot) performance across different models and languages. We use questions with answers followed by CoT prompt in the same language (native_cot) as the dataset and strict match score as the evaluation metric.
Appendix F Robustness Plots
F.1 CL-IFEval Plots
Figure 6: Cross-Lingual IFEval Aya-23-35B Model Comparison Across Languages
Figure 7: Cross-Lingual IFEval Gemma-2-9b-it Model Comparison Across Languages
Figure 8: Cross-Lingual IFEval Qwen3-8b Model Comparison Across Languages
F.2 CL-GSMSym Plots
Figure 9: Cross-Lingual GSMSym Aya-23-35B Template Comparison Across Languages for 50 generated samples of 10 templates
Figure 10: Cross-Lingual GSMSym Gemma-2-9b-it Template Comparison Across Languages for 50 generated samples of 10 templates
Figure 11: Cross-Lingual GSMSym Qwen3-8b Template Comparison Across Languages for 50 generated samples of 10 templates
Appendix G Failure Case Examples for CL-IFEval
[…]
Appendix H Template Examples for CL-GSMSym
To illustrate the structure and reasoning complexity of items in the CL-GSMSym benchmark, we present representative template-based examples
Template 3 Example (French) Language: French Template: 3 Question: Luis lance un dé à quatre faces. Quelle est la probabilité (exprimée en pourcentage) qu’il obtienne un nombre supérieur à 2 plutôt que deux nombres impairs consécutifs ? Answer: Il y a 2 nombres supérieurs à 2 sur le dé, donc les chances d’en obtenir un sont de $\frac{2}{4}=50\%$. La probabilité d’obtenir un nombre impair est de 50%, donc la probabilité d’en obtenir deux d’affilée est $0.5\times 0.5=25\%$. La différence entre ces deux probabilités est $50\%-25\%=25\%$. Final Answer : #### 25
Template 3 Example (English) Language: English Template: 3 Question: Faisal is rolling a four-sided die. How much more likely is it (expressed as a percentage) that he rolls a number greater than 1 than that he rolls two even numbers in a row? Answer: There are 3 numbers greater than 1 on the die, so the chance of rolling one is $\frac{3}{4}=75\%$. The chance of rolling one even number is 50%, so the chance of rolling two in a row is $0.5\times 0.5=25\%$. The difference between these probabilities is $75\%-25\%=50\%$. Final Answer: #### 50
Template 10 Example (Spanish) Language: Spanish Template: 10 Question: Cuando Emma observa a su primo, saca una variedad de juguetes para él. La bolsa de bloques de construcción tiene 74 bloques dentro. El contenedor de animales de peluche tiene 37 animales de peluche dentro. La torre de anillos apilables tiene 30 anillos multicolores. Emma compró recientemente un tubo de pelotas saltarinas, lo que eleva el número total de juguetes para su primo a 215. ¿Cuántas pelotas saltarinas vinieron en el tubo? Answer: Sea $T$ el número de pelotas que rebotan en el tubo. Después de comprar el tubo de pelotas, Emma tiene $21574+37+30+T=141+T=$ 215 juguetes. Por lo tanto, $T=215-141=74$. Final Answer: #### 74
Template 10 Example (English) Language: English Template: 10 Question: When Winnie watches her nephew, she gets out a variety of toys for him. The bag of building blocks has 72 blocks in it. The bin of stuffed animals has 47 stuffed animals inside. The tower of stacking rings has 29 multicolored rings on it. Winnie recently bought a tube of bouncy balls, bringing her total number of toys for her nephew up to 215. How many bouncy balls came in the tube? Answer: Let $T$ be the number of bouncy balls. Total number of toys before the balls: $72+47+29=148$. So $67148+T=215\Rightarrow T=215-148=$ 67. Final Answer: #### 67