### v2

# Multi-lingual Functional Evaluation for Large Language Models

Abstract
Multi-lingual competence in large language models is often evaluated via static data benchmarks such as Belebele, M-MMLU and M-GSM. However, these evaluations often fail to provide an adequate understanding of the practical performance and robustness of models across multi-lingual settings. In response, we create multi-lingual functional benchmarks – Cross-Lingual Grade School Math Symbolic (CL-GSM Symbolic) and Cross-Lingual Instruction-Following Eval (CL-IFEval)– by translating existing functional benchmark templates from English to five additional languages that span the range of resources available for NLP: French, Spanish, Hindi, Arabic and Yoruba. Our results show that the gap between static and functional evaluations is highly uneven: across models, performance drops from M-GSM to CL-GSM Symbolic by 24%, 17%, and 18% in English, French, and Spanish, while the drop from Belebele to CL-IFEval ranges from 15% to 24% across languages, and the drop from M-MMLU to CL-IFEval is much smaller (0.5% to 3%).Similarly, we find that model robustness across languages varies significantly, with certain languages (eg. Arabic, English) being the most consistently well performing across evaluation iterations.
Multi-lingual Functional Evaluation for Large Language Models
Victor Ojewale1,   Inioluwa Deborah Raji2,   Suresh Venkatasubramanian1 1The Center for Tech Responsibility, Brown University, USA 2University of California, Berkeley, USA
1 Introduction
Figure 1: Description of the functional evaluation paradigm. Unlike with static data benchmarks, in the functional evaluation paradigm, model input prompts are not fixed but generated through a fixed template and a set of variables $X$ (modifiable prompt attributes meant to impact model outputs) and a set of distractors $D$ (modifiable prompt attributes meant to be ignored). The ground truth in this setting is generated through a fixed functional transformation $f(X)$. For instance, the prompt "Sally bought 2 red apples and 3 green apples. How much fruit did Sally buy?" is generated from the fixed template $\{name\}$ bought $\{n_{1}\}$ $\{color_{1}\}$ apples and $\{n_{2}\}$ $\{color_{2}\}$ apples. How much fruit did $\{name\}$ buy?". This template involves the variables $X=\{n_{1},n_{2}\}$ and the distractors $D=\{name,color_{1},color_{2}\}$. The correct fixed output function in this case is $f(X)=n_{1}+n_{2}$.
Despite some meaningful progress, models operating in languages other than English have been regularly found to be more biased Talat et al. (2022), less safe Yong et al. (2023) and overall meaningfully less performant and robust  Ojo et al. (2023).
Popular multi-lingual LLM evaluations, such as Multilingual MMLU (M-MMLU) and the Multilingual Grade School Math (M-GSM) benchmark(Hendrycks et al., 2021; Shi et al., 2022; Cobbe et al., 2021), while useful, often fail to capture more meaningful indications of functional multi-lingual model performance – that is, the robust execution of a given prompt across a variety of languages (see Figure  1). In this paper, we extend the scope of two English "functional" evaluation datasets – IFEval Zhou et al. (2023), and GSM-Symbolic Mirzadeh et al. (2024) – by translating its prompt templates into five additional languages: French, Spanish, Hindi, Arabic and Yoruba.
Our experiments show that the relationship between functional and static benchmark performance varies significantly across models, languages, and benchmarks. While models often score higher on static benchmarks than on functional evaluations, the size of this gap varies substantially, with some static benchmarks aligning much more closely with functional performance than others. Language performance gaps also vary significantly across settings as their magnitude and direction shift across specific languages and models. Finally, functional benchmarks surface robustness that varies significantly by language and by prompt structure, revealing that models can be stable for certain languages yet brittle for others on particular instruction types, templates, or question families.
2 Related Work
Common multi-lingual data benchmarks such as M-MMLU (Hendrycks et al., 2021; Lai et al., 2023b), FLORES (Goyal et al., 2022), BeleBele (Bandarkar et al., 2024), and XLSum (Hasan et al., 2021) possess known limitations. The use of direct translation for many of these benchmarks has been critiqued by some as being devoid of realistic cultural context Romanou et al. (2024); Singh et al. (2024). Furthermore, research reveals that English language benchmark data contamination might distort reported benchmark performance in English or possibly additional languages (as is the case for MMLU Dodge et al. (2021) and GSM Zhang et al. (2024)).
Functional evaluation involves "templating" a common popular benchmark with modifiable variables. For example, GSM-Symbolic templates examples of the static data benchmark GSM8k  Cobbe et al. (2021) to generate the input permutations. The ground truth for the math problems is then calculated using literal template-based functional mappings from input values to the expected output (see Figure  1). Recent work has attempted to set up similar symbolic annotations for natural language benchmarks Hennigen et al. (2023), and we can see a similar verifiable, function-based template format with instruction-following benchmarks (Liu et al., 2024; Chang et al., 2023) such as the IFEval datasetZhou et al. (2023). Although some concurrent work – a proposed Multi-lingual IFEval (M-IFEval)  Dussolle et al. (2025) – has attempted to translate the IFEval template to Spanish, French and Japanese, we have yet to see more systematic analysis of how such functional evaluations can inform better assessments of performance and robustness in multi-lingual deployment settings.
Benchmark Avail. Lang Prompts/Lang Templates Samples/Template Task
Functional Benchmarks
Cross-Lingual GSM Symbolic (Ours) 5 5000 100 50 Math Reasoning
Cross-Lingual IFEval (Ours) 5 371 7 31 - 163 Instruction following
Static Benchmarks
Multilingual MMLU Lai et al. (2023b) 26 13062 N/A N/A MC knowledge recall
Multilingual GSM Shi et al. (2022) 10 250 N/A N/A Math Reasoning
Belebele Bandarkar et al. (2024) 122 900 N/A N/A MC reading comprehension
Table 1: Descriptive Statistics of Analyzed Benchmarks.
3 Benchmark Datasets and Templates
We construct multi-lingual functional benchmarks by translating (a) the 100 English GSM-Symbolic templates (Mirzadeh et al., 2024) and (b) the 541 English IFEval prompts (Zhou et al., 2023) each containing at least one verifiable instruction into French, Spanish, Yoruba, Hindi, and Arabic. Following Bang et al. (2023); Lai et al. (2023a), we refer to languages by ISO 639-1 codes and select them to span resource levels based on CommonCrawl coverage (Table 3).
Initial translations are produced with Google Translate (Wu et al., 2016). We then conduct native-speaker validation of all translated GSM-Symbolic templates and all translated IFEval prompts before any GSM-Symbolic item instantiation.
Validators and recruitment.
For each language, we recruited at least two native speakers volunteers to validate translations. Validators were recruited primarily from students studying at a US-based university. For Yoruba, we additionally recruited a native speaker based in Nigeria to ensure dialectal and cultural appropriateness in a low-resource setting.
Quality control, rubric, and prompt filtering.
Validation followed a structured rubric (Appendix G). Validators compared each translated prompt or template against the English source to ensure fidelity to meaning, intent, and constraint strength, and to verify that instruction realizations remained valid in the target language. Validators discussed and, when possible, corrected problematic translations, and also checked to detect misspellings, duplicated sentences, and template–instance alignment issues. Validators noted that occasional literal word choices remained but did not alter functional constraints or ground-truth mappings.
As part of this validation, they also flagged a subset of original IFEval prompts whose instructions are not meaningfully evaluable in scripts without case distinction (e.g., uppercase/lowercase constraints in Arabic and Hindi). We removed these prompts to preserve cross-lingual comparability and evaluation validity, yielding a 371-prompt CL-IFEval set.
After validation, each GSM-Symbolic template was instantiated into 50 variants per language, yielding 5,000 QA pairs per language, while the validated IFEval prompts constitute CL-IFEval.
For comparison with multi-lingual static benchmarks, we evaluate on M-MMLU (Lai et al., 2023b; Hendrycks et al., 2021), M-GSM (Shi et al., 2022), and Belebele (Bandarkar et al., 2024). The Cross-Lingual GSM Symbolic and Cross-Lingual IFEval datasets are publicly available.  [^1] https://huggingface.co/datasets/vojewale/Cross-lingualGSMSymbolic [^2] https://huggingface.co/datasets/vojewale/Cross-lingualIFEval
Further dataset details can be found in Table  1.
4 Model Evaluation
For tasks that benefit from step-by-step reasoning, we use Chain-of-Thought (CoT) prompting in the same language as the query where applicable. Unless otherwise stated, CL-GSMSym follows the native 8-shot GSM8K setting, and MGSM uses the native_cot configuration. All evaluations are orchestrated with the EleutherAI LM Evaluation Harness (lm-eval), which standardizes prompts, decoding, and metrics and is widely used for open-weight models (Gao et al., 2024).
For reproducible consistency in reporting, our evaluations spanned open-source large language models, including instruction-tuned and multilingual variants—Aya 23-35B, Aya Expanse-32B, Gemma-2-9B-it, Qwen3-8B, Mistral-7B-Instruct-v0.3, and Mixtral-8x7B-Instruct-v0.1. Using CL-IFEval, we compute prompt-level strict/loose and instruction-level strict/loose accuracy as in Zhou et al. (2023). For CL-GSMSym, after template translation, we instantiate 50 variants per template per language and evaluate a random sample of 500 items per language under the native 8-shot GSM8K setting.
Refer to captionRefer to captionRefer to caption
Figure 2: Correlation plot of Performance Gap between MGSM, MMLU, Belebele (left-to-right) and and CL-IFEval for High Resourced Languages only (en, fr, es). This reveals that measured language performance gaps (i.e. the difference between the performance on the highest performant language and the lowest performant language) are notably larger in functional evaluations than in static data benchmarks.
5 Results
5.1 Average Performance Across Languages & Model Ranking
Models perform consistently high for high-resourced languages on static natural language benchmarks (eg. For M-MMLU, English (66% - 69%), French (51% - 65%) and Spanish (51% - 65%) performance is consistently higher; for Belebele, performance across models for high-resource languages is even stronger (79% – 90%)). However, there is notable performance variation across models for mathematical reasoning as measured with static data benchmarks – for example, M-GSM English scores range from 44% to 86% across models, while performance in French and Spanish is generally lower, though still relatively competitive and wide-ranging (French: 44% - 76%, Spanish: 37% - 84%). When available, performance across models decreases on medium resource languages - with scores ranging from 31% to 48% in Hindi for M-MMLU, and 41% to 74% in Hindi for Belebele. Low-resource languages like Yoruba show drastic performance drops, with scores below 32% on Belebele.
Functional benchmarks expose larger within-language variability.
On the other hand, functional benchmarks in natural language, such as CL-IFEval, show a wide range on even English performance (56% - 88%), with other high and medium resource languages such as French (36%- 62%), Spanish (33% - 57%), and Arabic (20% - 50%) being similarly wide ranging in performance. CL-GSMSym shows a similar pattern, with a wide range of model performance in high and medium resource languages (English: 49% - 87%; French: 39% - 78%; Spanish: 34% - 82%; Arabic: 18% - 79%; Hindi: 15% - 72%). Interestingly, in both cases, performance for low resource languages was more consistently low (e.g. in Yoruba, performance on CL-IFEval ranges from 12% to 23%, and performance on CL-GSMSym ranges from 4% to 14%).
Model tiering and ranking shifts under functional evaluation.
For both static benchmarks, the same models are consistently the most performant - typically Qwen3-8B, followed by either Gemma-2-9b-it or Aya-expanse-32B - with the other models not far behind. For functional benchmarks there is a clear discrepancy between a performant class of models (Qwen3-8B, followed by Aya-expanse-32B, followed by Gemma-2-9b-it) and the rest of the models. Unlike with static data benchmarks, this stark performance discrepancy persists, even for low resource languages. For example, on CL-IFEval, the top three models have an English performance range of 70% to 88%, whereas the bottom three models all have an English performance range consistently around 60%. In Yoruba, this performance class discrepancy is still observed, though the performance gap between models is much more narrow. Comparatively, the top three models on the static data benchmark Belebele perform in English at 90% to 93% and the bottom three models perform at 80% to 86%. Across languages, we can see model rankings change significantly when assessed with static vs functional benchmarking. For example, a typically lower-ranked model like Aya-23-35B, outranks Aya-expanse-32B and Qwen3-8B in Yoruba on the static Belebele benchmark but not in Yoruba on the functional CL-IFEval benchmark. Similarly, the Aya-23-35B model outranks Gemma-2-9b-it generally on the static M-GSM, even though it consistently performs worse than Gemma-2-9b-it on the functional CL-GSMSym benchmark, for the same languages.
Full details of the performance results for the static data benchmarks and functional benchmarks can be found in Appendix D and E.
5.2 Language Performance Gap
We define the language performance gap to be the difference of the model’s accuracy on its lowest performant language, and the highest performant language. For high- resourced languages, functional evaluations tend to reveal a much larger performance discrepancy than static data benchmarks (see Figure  2). Notably, Aya-expanse-32B, a Cohere model marketed specifically for its multi-lingual capability, has a 5.89% average error gap between high-resourced languages (en, es, fr) across static data benchmarks, but a 25.88% performance gap between those same languages on the functional benchmark CL-IFEval. Similarly, Qwen3-8b has fairly low language performance gaps across high resourced languages on static data benchmarks like M-GSM (14.4%), M-MMLU (4.78%), and Belebele (3.00%), but a large language performance gap on the functional CL-IFEval benchmark (31%). Interestingly, for mathematical reasoning, the opposite can be true – for example, the language performance gap across high resource languages is lower for the functional CL-GSMSym (e.g. for Qwern3-8b, 8.6%) than the static data benchmark M-GSM (eg. for Qwern3-8b, 14.4%). Similarly, when including low resource language settings, CL-IFEval tends to show more optimistic numbers (a higher average performance, a lower performance gap) than the static data benchmark, Belebele. In Appendix B, Tables 4, 5,6, as well as Figure  2, we provide further details on these results.
5.3 Instruction and Template-Level Robustness Across Languages
Instruction-group breakdowns reveal where cross-lingual generalization fails.
To evaluate instruction-level robustness across languages, we use CL-IFEval to compare model behavior on the IFEvalZhou et al. (2023) default grouped instruction categories. Figures 3, 4, and 5 show success rates for Aya-23-35B, Gemma-2-9B, and Qwen3-8B respectively. These plots reveal notable variation in cross-lingual generalization, with Yoruba performing the worst across most instruction groups and showing complete failure in the “Start/End” category for all models.
Template-controlled math variants isolate which problem families drive errors.
To further probe model robustness under controlled prompt variations, we use a subset of CL-GSMSym consisting of 50 samples generated from a fixed set of 10 mathematical question templates from the GSM Symbolic datasetMirzadeh et al. (2024). Figures 6, 7, and 8 display the per-template accuracy of the same three models across languages. Template 3(Appendix I), which involves probabilistic inference, consistently produced the lowest performance across models and languages highlighting a persistent weakness in generalizing probabilistic reasoning across linguistic contexts. Interestingly, performance in certain middle-resourced languages (ar, hi) is just as if not more robust for certain types of problems than high resourced languages (e.g. in CL-GSM Symbolic, see Template 2, 4 and 10 results in Figure  6).
6 Conclusion
We introduce two new benchmarks, CL-IFEval and CL-GSMSym, for multi-lingual functional evaluation. Our experiments uncovered major language performance gaps across languages, even for LLMs with robust multilingual claims and strong static data benchmark scores.
Limitations.
We use automated translation tools like Google Translate in constructing CL-IFEval and CL-GSMSym. While these tools offer broad language coverage and facilitate large-scale data generation, they introduce potential inaccuracies, particularly for lower resourced languages like Yoruba, and when dealing with conversions across metric and imperial measurement systems.
Another limitation is the emphasis in our analysis on open-weight models. Our primary focus remains on open-weight models such as Mixtral-8x7B, Mistral-7B, Gemma2-9B-it, Qwen3-8B and AYA models. This creates a possible inherent bias in the scope of our comparisons, as proprietary models may perform significantly better than their open-weight counter-parts. On the other hand, the consistency and transparency of open-weight models make them the preferable object of study – the incorporation of proprietary models can make results hard to reproduce reliably.
Future Work.
Future directions include systematically curating higher-quality translations, expanding into multi-lingual code or multi-modal instruction evaluation, and further investigating the robustness and error patterns in both functional and static benchmarks. Also, as functional evaluations involve automatic verification, there is some possibility of extrapolating this framework in the training and evaluation of multi-lingual reasoning models Yong et al. (2025).
Acknowledgments
The authors would like to thank Zheng-Xin Yong for feedback on the work. This work was supported in part by the MacArthur Foundation, the Mozilla Foundation, and the Heising-Simons Foundation.
Appendix A Additional Experimental Details
Benchmark Avail. Lang Prompts/Lang Templates Samples/Template Task
Functional Benchmarks
Cross-Lingual GSM Symbolic (Ours) 5 5000 100 50 Math Reasoning
Cross-Lingual IFEval (Ours) 5 371 7 31 - 163 Instruction following
Static Benchmarks
Multilingual MMLU Lai et al. (2023b) 26 13062 N/A N/A MC knowledge recall
Multilingual GSM Shi et al. (2022) 10 250 N/A N/A Math Reasoning
Belebele Bandarkar et al. (2024) 122 900 N/A N/A MC reading comprehension
Table 2: Descriptive Statistics of Analyzed Benchmarks.
A.1 Models
Our evaluations spanned open-source large language models, including instruction-tuned and multilingual variants. Below, we provide details on each model:
  • Aya 23-35B: A multilingual instruction-tuned model based on Cohere’s Command framework (Aryabumi et al., 2024).
  • Aya Expanse-32B: Part of the Aya Expanse series, designed to enhance multilingual performance capabilities(Dang et al., 2024).
  • Gemma-2-9B-it: An instruction-tuned model trained on 8 trillion tokens from diverse sources, including web documents, code, and mathematical text. It is optimized for a wide range of text generation tasks, including question answering, summarization, and code understanding (Team et al., 2024).
  • Qwen3-8B: A dense and mixture-of-experts (MoE) model supporting 100+ languages. It includes specialised modes for logical reasoning, code generation, and agent-based tasks (Yang et al., 2025).
  • Mistral-7B-Instruct-v0.3: An open-source instruction fine-tuned variant of Mistral-7B (Jiang et al., 2023).
  • Mixtral-8x7B-Instruct-v0.1: (Jiang et al., 2024)A sparse mixture-of-experts model with 12.9B active parameters per token, fine-tuned for instruction tasks.
A.2 Language Resource Classification
Table 3 classifies the languages used in our evaluation based on their relative resource levels in CommonCrawl.
Language CC-MAIN-2025-05 (%) Category
English (en) 43.37 High-resource (HRL)
French (fr) 4.52 High-resource (HRL)
Spanish (es) 4.64 High-resource (HRL)
Arabic (ar) 0.66 Medium-resource (MRL)
Hindi (hi) 0.20 Medium-resource (MRL)
Yoruba (yo) 0.0009 Extremely low-resource (X-LRL)
Table 3: Classification of languages based on resource availability in CommonCrawl ($C$-MAIN-2025-05).
Appendix B Full Performance Gap Results
B.1 High-Resource Language Performance Gaps
Model CL-IFEval CL-GSMSym M-GSM M-MMLU Belebele
Aya-23-35B 26.68 4.4 8.0 4.63 5.00
Mixtral-8x7B-Instruct 22.36 7.0 12.0 3.63 6.89
Mistral-7B-Instruct 18.60 14.0 8.0 7.69 15.88
Aya-expanse-32B 25.88 6.2 9.6 3.84 4.22
Gemma-2-9b-it 24.80 27.4 29.2 6.39 3.66
Qwen3-8b 31.00 8.6 14.4 4.78 3.00
Table 4: Performance Gap on High-Resource Languages, as measured by the benchmark accuracy difference between the results on the best performant language and worst performant language across the set of analyzed languages (en, fr, es).
B.2 High to Medium-Resource Language Performance Gaps
Model CL-IFEval CL-GSMSym M-MMLU Belebele
Aya-23-35B 26.68 17.6 18.08 22.45
Mixtral-8x7B-Instruct 31.41 33.2 30.65 36.78
Mistral-7B-Instruct 37.74 33.8 27.92 38.44
Aya-expanse-32B 28.03 11.0 15.29 15.45
Gemma-2-9b-it 28.04 27.4 18.20 19.55
Qwen3-8b 37.73 27.4 20.03 23.45
Table 5: Performance Gap on High to Medium-Resource Languages, as measured by the benchmark accuracy difference between the results on the best performant language and worst performant language across the set of analyzed languages (en, fr, es, ar, hi).
B.3 High to X Low-Resource Language Gaps
Model CL-IFEval CL-GSMSym Belebele
Aya-23-35B 51.21 57.4 54.12
Mixtral-8x7B-Instruct 43.66 56.8 57.11
Mistral-7B-Instruct 40.98 44.2 50.11
Aya-expanse-32B 56.87 71.6 60.45
Gemma-2-9b-it 52.29 64.4 52.44
Qwen3-8b 69.81 73.0 62.89
Table 6: Performance Gap on High to X Low-Resource Languages, as measured by the benchmark accuracy difference between the results on the best performant language and worst performant language across the set of analyzed languages (en, fr, es, ar, hi, yo).
Appendix C Static Data Benchmark Results
C.1 Multilingual MMLU (M-MMLU)
The M-MMLU benchmark assesses knowledge and reasoning in a wide range of subjects and languages. The results in Table 11 demonstrate that performance is generally higher in English (0.66–0.69) and other well-resourced languages such as French (0.51–0.65) and Spanish (0.51–0.65). In contrast, performance in medium-resource languages like Hindi is considerably lower, with scores ranging from 0.31 to 0.48. Arabic performance falls between these extremes but is still notably below English and French levels.
C.2 Multilingual Grade School Math benchmark (MGSM)
MGSM tests a model’s capacity to handle multi-step reasoning problems, particularly in mathematical contexts. The results in Table 12 reveal a stark performance gap between English and other languages. English scores range from 0.44 to 0.86, while performance in French and Spanish is generally lower, though still relatively competitive (French: 0.44–0.76, Spanish: 0.37–0.84). French and Spanish scores perform better than seen in M-MMLU, but medium and low-resource languages are not included in this benchmark.
C.3 Belebele Evaluation
The Belebele benchmark Bandarkar et al. (2024) assesses reading comprehension by requiring models to answer multiple-choice questions across a wide variety of languages. Table 10 highlights significant disparities in performance across resource levels. English comprehension is consistently strong, with scores around 0.85–0.90, while other high-resource languages such as French and Spanish achieve similarly high performance (0.79–0.88). However, low-resource languages like Yoruba show drastic performance drops, with scores below 0.32. Hindi also exhibits relatively weak results (0.41–0.74).
Appendix D Full Functional Benchmark Results
Model English French Spanish Arabic Hindi Yoruba
AYA Exp-32B 74.93 55.52 49.05 46.90 49.86 18.06
AYA 23-35B 67.11 48.78 40.43 43.93 40.70 15.90
Mistral-7B-Instruct 59.57 41.78 40.97 25.33 21.83 18.59
Mixtral-8x7B-Instruct 56.06 36.39 33.70 31.53 24.65 12.40
Gemma-2-9B-it 75.20 53.36 50.40 47.16 48.24 22.91
Qwen3-8B 87.60 61.46 56.60 50.13 49.87 17.79
Table 7: Cross-Lingual-IFEval Prompt-level Strict performance across different models and languages.
Language Model Strict PL (%) Strict IL (%) Loose PL (%) Loose IL (%)
English AYA 23-35B 67.11 73.75 70.89 77.41
Gemma2-9B-it 75.20 81.08 79.24 84.17
AYA Exp-32B 74.93 79.92 78.97 83.39
Mistral-7B 59.57 68.15 64.42 72.39
Mixtral-8x7B 56.06 64.48 61.19 69.50
Qwen3-8B 87.60 90.70 91.37 93.63
French AYA 23-35B 48.78 58.30 53.90 62.94
Gemma2-9B-it 53.36 61.38 56.06 63.89
AYA Exp-32B 55.52 63.12 58.22 65.83
Mistral-7B 41.78 51.35 46.90 56.18
Mixtral-8x7B 36.39 45.37 39.62 49.03
Qwen3-8B 61.46 69.11 63.07 70.46
Spanish AYA 23-35B 40.43 50.77 44.20 54.24
Gemma2-9B-it 50.40 59.07 51.75 61.00
AYA Exp-32B 49.05 58.49 51.21 60.62
Mistral-7B 40.97 50.96 43.66 53.86
Mixtral-8x7B 33.70 43.47 37.20 47.00
Qwen3-8B 56.60 65.06 57.41 65.83
Yoruba AYA 23-35B 15.90 22.77 16.17 23.74
Gemma2-9B-it 22.91 31.08 25.88 34.75
AYA Exp-32B 18.06 28.19 29.48 30.31
Mistral-7B 18.59 28.95 20.21 31.27
Mixtral-8x7B 12.40 20.27 15.00 23.36
Qwen3-8B 17.79 27.22 20.21 30.50
Arabic AYA 23-35B 43.93 53.47 46.90 56.37
Gemma2-9B-it 47.16 57.14 48.79 58.88
AYA Exp-32B 46.90 56.76 49.06 58.69
Mistral-7B 25.33 36.10 28.57 39.38
Mixtral-8x7B 31.53 40.73 35.85 45.95
Qwen3-8B 50.13 60.62 52.30 62.16
Hindi AYA 23-35B 40.70 50.19 44.47 54.44
Gemma2-9B-it 48.24 56.94 50.94 59.65
AYA Exp-32B 49.86 58.69 52.56 61.78
Mistral-7B 21.83 32.24 25.06 36.67
Mixtral-8x7B 24.65 33.24 32.83 41.31
Qwen3-8B 49.87 58.11 52.02 60.62
Table 8: Comprehensive instruction-following results across multiple models and languages. PL = Prompt-Level Accuracy; IL = Instruction-Level Accuracy.
Model English French Spanish Arabic Hindi Yoruba
Aya-23-35B 56.00 55.00 59.40 61.20 43.60 3.80
Mixtral-8x7B-Instruct 56.80 55.00 62.00 32.40 28.80 5.20
Mistral-7B-Instruct 48.60 39.40 34.60 18.40 14.80 4.40
Aya-expanse-32B 82.60 76.40 82.20 76.80 71.60 11.00
Gemma-2-9b-it 77.00 49.60 54.80 55.20 57.00 12.60
Qwen3-8B 86.80 78.20 81.80 78.80 59.40 13.80
Table 9: CL-GSMSym (8-shot) performance across different models and languages.
Appendix E Full Static Data Benchmark Results
Model English French Spanish Arabic Hindi Yoruba
Aya-23-35B 0.8556 0.8133 0.8056 0.7667 0.6311 0.3144
Mixtral-8x7B-Instruct 0.8522 0.8100 0.7833 0.6344 0.4844 0.2811
Mistral-7B-Instruct 0.7944 0.6711 0.6356 0.5122 0.4100 0.2933
Aya-expanse-32B 0.8989 0.8844 0.8567 0.8367 0.7444 0.2944
Gemma-2-9b-it 0.9333 0.92 0.8967 0.8867 0.7378 0.4089
Qwen3-8B 0.9256 0.9089 0.8956 0.87 0.6911 0.2967
Table 10: Belebele performance across different models and languages.
Model English French Spanish Arabic Hindi
Aya-23-35B 0.6617 0.6168 0.6154 0.5394 0.4809
Mixtral-8x7B-Instruct 0.6854 0.6508 0.6491 0.4142 0.3789
Mistral-7B-Instruct 0.5900 0.5131 0.5181 0.3328 0.3108
Aya-expanse-32B 0.7396 0.7012 0.7025 0.6267 0.5867
Gemma-2-9b-it 0.7149 0.6635 0.6541 0.5565 0.5329
Qwen3-8B 0.7528 0.7050 0.7091 0.6072 0.5525
Table 11: M-MMLU(5-shot) performance across different models and languages.
Model English French Spanish
Aya-23-35B 0.668 0.588 0.600
Mixtral-8x7B-Instruct 0.632 0.512 0.536
Mistral-7B-Instruct 0.448 0.444 0.368
Aya-expanse-32B 0.856 0.760 0.840
Gemma-2-9b-it 0.664 0.372 0.512
Qwen3-8B 0.880 0.736 0.852
GPT-4o-mini 0.656 0.556 0.432
Claude 3.5 Sonnet 0.900 0.716 0.812
Table 12: MGSM (5-shot) performance across different models and languages. We use questions with answers followed by CoT prompt in the same language (native_cot) as the dataset and strict match score as the evaluation metric.
Appendix F Robustness Plots
F.1 CL-IFEval Plots
Figure 3: Cross-Lingual IFEval Aya-23-35B Model Comparison Across Languages
Figure 4: Cross-Lingual IFEval Gemma-2-9b-it Model Comparison Across Languages
Figure 5: Cross-Lingual IFEval Qwen3-8b Model Comparison Across Languages
F.2 CL-GSMSym Plots
Figure 6: Cross-Lingual GSMSym Aya-23-35B Template Comparison Across Languages for 50 generated samples of 10 templates
Figure 7: Cross-Lingual GSMSym Gemma-2-9b-it Template Comparison Across Languages for 50 generated samples of 10 templates
Figure 8: Cross-Lingual GSMSym Qwen3-8b Template Comparison Across Languages for 50 generated samples of 10 templates
Appendix G Translation Validation Rubrics
G.1 CL-IFEval Translation Validation (541 Prompts)
1. Objective The goal of this validation task is to ensure that every translated prompt accurately represents the intent, structure, and constraints of the original English version while remaining clear and natural in the target language. Annotators should: Confirm that the translation faithfully conveys the same meaning and instructions as the English source. Verify that every element in the instruction_id_list is valid and meaningful in the target language. Flag any instruction IDs that cannot be reasonably expressed or understood in that language. Correct minor translation errors or omissions directly in the prompt, but only when needed for consistency with the English version. Record every correction or flag for traceability.
2. Dataset Structure Each entry contains: key – unique identifier for the instance prompt – translated text (the only part that differs from English) instruction_id_list – the list of instructions applied to the English prompt kwargs – parameters or values associated with those instructions Example (simplified):
"key": 1001,
"prompt": "Translated prompt text...",
"instruction_id_list": ["punctuation:no_comma", "length_constraints:number_words"],
"kwargs": [{}, {"num_words": 300}]
3. Validation Criteria A. Translation Fidelity The translation must convey the same meaning, intent, and constraints as the English prompt. The tone and phrasing should match the formality and directive style of the English source. If the translation is missing, ambiguous, or distorts the meaning, revise the prompt to align with the English intent. When a fix is made, record it with a short note (e.g., “clarified no-comma rule,” “added minimum-count phrase”). Do not add new content or stylistic embellishments not present in the English prompt. B. Instruction–Prompt Alignment For every instruction in the instruction_id_list: * Realized – The translated prompt clearly expresses the instruction. * Not Applicable – The instruction refers to a feature or rule that cannot exist in the target language (e.g., case sensitivity in scripts that lack uppercase/lowercase). Keep the ID in the metadata, but flag it as “not applicable in this language.” Do not remove or modify IDs in the JSON. All flags are recorded separately. C. Linguistic and Logical Consistency The translation must be grammatically correct, culturally neutral, and logically coherent. Instructions like “Do not use commas” or “Write at least 300 words” must retain their restrictive force and avoid weaker formulations such as “Try not to use.” Quantitative and conditional phrases (e.g., “at least,” “exactly,” “no more than”) must match the English structure precisely. D. Formatting and Style Preserve examples, symbols, and URLs exactly as they appear. Maintain Markdown examples (*highlighted section*, headings, etc.) where relevant. Ensure punctuation, numbering, and quotation marks follow conventions appropriate for the target language. Do not introduce additional formatting not found in the original.
4. Correction and Flagging Process For each entry: 1. Compare the English prompt with the translated version. 2. Identify whether each instruction ID is not applicable. 3. If small translation inconsistencies exist, correct them directly in the prompt. 4. Record every modification or flag, even for minor fixes. 5. Do not introduce new instructions or stylistic changes that alter the intent. Permitted corrections include: Fixing missing or mistranslated phrases that change meaning. Clarifying numerical constraints or prohibitions (e.g., “Do not use commas”). Adjusting phrasing for grammatical correctness. Not permitted: Adding new examples, stylistic flourishes, or re-interpreting the prompt. Changing the number of constraints or rewriting the task.
5. Final Review and Consistency Check Before final submission: Ensure that all translations are accurate and natural in their respective languages. Verify that English-only or script-dependent instruction IDs (e.g., case-based rules) are consistently flagged as “Not Applicable.” Confirm that all modifications and flags have been logged. Export the validated, corrected dataset for each language along with its annotation log in the Log changes column.
G.2 CL-GSMSym Translations Validation Guidelines (100 Math Templates)
1. Objective The goal of this validation task is to ensure that every translated mathematical template accurately preserves the structure, logic, and functionality of the original English version. Annotators should verify that translations remain technically valid, mathematically consistent, and compatible with the generator framework. The task involves: Preserving all template markup ({…}, #init:, #conditions:, #answer:). Ensuring the math logic and units match the original. Correcting small formatting and spacing issues where necessary. Flagging every fix or inconsistency for audit tracking.
2. Dataset Components Each template file contains: question – the main text of the math problem. answer – the step-by-step explanation and final numeric answer. question_annotated – question with placeholders, variables, and meta blocks. answer_annotated – annotated solution including variables and computed forms. #init, #conditions, #answer – meta-sections used by the template generator. id_orig / id_shuffled – unique identifiers for the template.
3. Validation Criteria A. Template Markup Preservation Every variable token must remain identical to English: {var}, {var,label}, {x,10}, {k*y}, etc. The meta-sections (#init:, #conditions:, #answer:) must remain in English and unaltered. If a translation changed a placeholder (for example, {big_fish,shark} $\rightarrow$ {big_fish,requin}), restore the English version. Flag: “Restored placeholder token”. B. Math and Unit Consistency All numeric quantities, operations, and measurement units must match the logic in the English template. Example: if English uses “6 inches”, “10-foot shark”, “12 inches/foot”, the translation must also use imperial units, not metric. If the translation converted units (e.g. inches $\rightarrow$ cm, feet $\rightarrow$ meters), revert to the English units. Flag: “Unit drift fixed (cm $\rightarrow$ inches)”. C. Angle Brackets and Final Answer Maintain the <<...>> spans exactly as in the English version (these are computation traces). The final line of every answer must end with: #### <number> or #### {expression}. If extra spaces, broken symbols, or HTML entities appear, normalize and flag. Flag: “Computation span or final answer normalized”. D. Meta Sections (Technical Integrity) The #init:, #conditions:, and #answer: sections must remain in English and in the same order. If the translation system or annotator translated these lines, restore them verbatim from the original English. Flag: “Meta block reverted to EN”. Ensure indentation and syntax are unchanged — the generator relies on exact structure.
4. Correction and Flagging Process For each template: 1. Open both the English and translated versions. 2. Compare them line by line (especially question_annotated and answer_annotated). 3. Apply small corrections directly: Restore missing or broken placeholders. Fix spacing and punctuation. Revert units if changed. 4. Record every modification you make, even small ones. 5. If the template cannot be fixed to run logically, flag it for exclusion.
5. Final Review Checklist Before final submission, verify that: All placeholders {...} and computation spans <<...>> are intact. All templates end with the correct final answer format. Meta sections remain in English, properly indented. All numbers and units align with the original logic. Each modification has been recorded in your correction log.
6. Output The deliverables per language: A clean, generator-compatible set of 100 validated templates. A Column in the google sheet with every correction or flagged issue (for transparency and reproducibility).
Appendix H Failure Case Examples for CL-IFEval
Failure Example: length_constraints:number_words Model: Qwen3-8B Language: English Prompt: Create a riddle about the name Sheldon using only 10 words. Make sure to only use capital letters in your entire response. Model Response: SHELDON, A NAME WITH A FUNKY RING. Instructions: change_case:english_capital length_constraints:number_words Failure: The response contained only 8 words, violating the 10-word constraint.
Failure Example: change_case:capital_word_frequency Model: Qwen3-8B Language: French Prompt: Rédigez un CV pour Matthias Algiers. Utilisez des mots en majuscules pour mettre en avant les compétences clés, mais assurez-vous que les mots entièrement en majuscules apparaissent moins de 10 fois. Enveloppez toute la réponse entre guillemets. Model Response: "MATTHIAS ALGIERS RÉSUMÉ PROFESSIONNEL Professionnel dévoué et orienté résultats avec une solide expérience en GESTION DE PROJET, PLANIFICATION STRATÉGIQUE et LEADERSHIP D’ÉQUIPE…" Instructions: change_case:capital_word_frequency startend:quotation Failure: The response included more than 10 fully capitalized words, violating the capital word frequency constraint.
Appendix I Template Examples for CL-GSMSym
To illustrate the structure and reasoning complexity of items in the CL-GSMSym benchmark, we present representative template-based examples
Template 3 Example (French) Language: French Template: 3 Question: Luis lance un dé à quatre faces. Quelle est la probabilité (exprimée en pourcentage) qu’il obtienne un nombre supérieur à 2 plutôt que deux nombres impairs consécutifs ? Answer: Il y a 2 nombres supérieurs à 2 sur le dé, donc les chances d’en obtenir un sont de $\frac{2}{4}=50\%$. La probabilité d’obtenir un nombre impair est de 50%, donc la probabilité d’en obtenir deux d’affilée est $0.5\times 0.5=25\%$. La différence entre ces deux probabilités est $50\%-25\%=25\%$. Final Answer : #### 25
Template 3 Example (English) Language: English Template: 3 Question: Faisal is rolling a four-sided die. How much more likely is it (expressed as a percentage) that he rolls a number greater than 1 than that he rolls two even numbers in a row? Answer: There are 3 numbers greater than 1 on the die, so the chance of rolling one is $\frac{3}{4}=75\%$. The chance of rolling one even number is 50%, so the chance of rolling two in a row is $0.5\times 0.5=25\%$. The difference between these probabilities is $75\%-25\%=50\%$. Final Answer: #### 50
Template 10 Example (Spanish) Language: Spanish Template: 10 Question: Cuando Emma observa a su primo, saca una variedad de juguetes para él. La bolsa de bloques de construcción tiene 74 bloques dentro. El contenedor de animales de peluche tiene 37 animales de peluche dentro. La torre de anillos apilables tiene 30 anillos multicolores. Emma compró recientemente un tubo de pelotas saltarinas, lo que eleva el número total de juguetes para su primo a 215. ¿Cuántas pelotas saltarinas vinieron en el tubo? Answer: Sea $T$ el número de pelotas que rebotan en el tubo. Después de comprar el tubo de pelotas, Emma tiene $74+37+30+T=141+T=215$ juguetes. Por lo tanto, T=215141=74T=215-141=74. Final Answer: #### 74
Template 10 Example (English) Language: English Template: 10 Question: When Winnie watches her nephew, she gets out a variety of toys for him. The bag of building blocks has 72 blocks in it. The bin of stuffed animals has 47 stuffed animals inside. The tower of stacking rings has 29 multicolored rings on it. Winnie recently bought a tube of bouncy balls, bringing her total number of toys for her nephew up to 215. How many bouncy balls came in the tube? Answer: Let $T$ be the number of bouncy balls. Total number of toys before the balls: $72+47+29=148$. So $148+T=215\Rightarrow T=215-148=67$. Final Answer: #### 67